Angel is twice as old as Gavin. In 3 years, Angel will be 13 years less than twice as old as Gavin. Find their ages now.
Answer by HyperBrain(694)

Let x=Angel's present age
y=Gavin's present age
x=2y
x+3=2(y+3)-13
Substituting x=2y,
2y+3=2y+6-13=2y-7
3=-7----false!
\
Therefore, there is no solution.
Power up,
HyperBrain!
Stop this mess! Click here